OpenStudy (anonymous):

Solve the equation 7 + n + 1 = 5(1- n) for n.

jigglypuff314 (jigglypuff314):

you understand how to distribute?

OpenStudy (anonymous):

Some what, would you explain? @jigglypuff314

OpenStudy (anonymous):

7+n+1=5-5n n+5n=-7-1+5 n=-1/2

jigglypuff314 (jigglypuff314):

so if a(x+y) then distributing means ax + ay so if 5(1-n) then 5(1) - 5(n)

OpenStudy (anonymous):

So the answer would be 0.5? or 1.5?

OpenStudy (anonymous):

Thanks guys! :)

jigglypuff314 (jigglypuff314):

neither... 7 + n + 1 = 5(1- n) so you would first distribute so 7 + n + 1 = 5(1) - 5(n) then simplify n + 8 = 5 - 5n then add 5n to both sides +5n +5n so 6n + 8 = 5 - 0 then subtract 8 on both sides -8 -8 6n = -3 then divide both sides by 6 /6 /6 so n = -3/6 = -1/2 = -0.5